# RAttention: Towards the Minimal Sliding Window Size in Local-Global Attention Models

### v1

@@ Abstract
Local-global attention models characterAI2024 ; gemmateam2024gemma2 have recently emerged as compelling alternatives to standard Transformers, promising improvements in both training and inference efficiency. However, the crucial choice of window size presents a Pareto tradeoff: larger windows maintain performance akin to full attention but offer minimal efficiency gains in short-context scenarios, while smaller windows can lead to performance degradation. Current models, such as Gemma2 and Mistral, adopt conservative window sizes (e.g., 4096 out of an 8192 pretraining length) to preserve performance. This work investigates strategies to shift this Pareto frontier, enabling local-global models to achieve efficiency gains even in short-context regimes. Our core motivation is to address the intrinsic limitation of local attention—its complete disregard for tokens outside the defined window. We explore RAttention, a variant of local attention integrated with a specialized linear attention mechanism designed to capture information from these out-of-window tokens. Pretraining experiments at the 3B and 12B scales demonstrate that RAttention achieves a superior Pareto tradeoff between performance and efficiency. As a sweetspot, RAttention with a window size of just 512 consistently matches the performance of full-attention models across diverse settings. Furthermore, the recurrent nature inherent in the linear attention component of RAttention contributes to enhanced long-context performance, as validated on the RULER benchmark hsieh2024ruler . Crucially, these improvements do not compromise training efficiency; thanks to a specialized kernel implementation and the reduced window size, RAttention maintains training speeds comparable to existing state-of-the-art approaches.  [^1] Our JAX implementation of RAttention using Pallas kernels will be open-sourced at https://github.com/apple/axlearn/tree/main/axlearn/common/rattention after internal review.
@@ 3.1 Residual Linear Attention (RLA)
where instead of reading out from ${\mathbf{S}}_{t}$ as in Eq 1, ${\bm{o}}_{t}$ is obtained via reading out from ${\mathbf{S}}_{t-w-1}$ – the hidden states that captures the contextual information ends at token $(t-w-1)$. To integrate with SWA, we combine he outputs from RLA and SWA, as demonstrated in Figure 1.
@@ 3.2 Hybrid RAttention Models
inter-chunkinter-chunkintra-chunkintra-chunk $\mathbf{S}_{[i-1]}$ ${\mathbf{S}}_{[i]}$ldots ${\mathbf{S}}_{[\scriptscriptstyle i+m-1]}$ ${\mathbf{S}}_{[i+m]}$ ${\mathbf{O}}^{\text{inter}}_{[i]}$ ${\mathbf{O}}^{\text{intra}}_{[i]}$ ${\mathbf{O}}^{\text{inter}}_{[i+m]}$ ${\mathbf{O}}^{\text{intra}}_{[i+m]}$ ${\mathbf{Q}}_{[i]}$ ${\mathbf{K}}_{[i]}$ ${\mathbf{V}}_{[i]}$ ${\mathbf{Q}}_{[\scriptscriptstyle i+m]}$ ${\mathbf{K}}_{[\scriptscriptstyle i+m]}$ ${\mathbf{V}}_{[\scriptscriptstyle i+m]}$ Load from HBMStore to HBMOn-chipxx Sequential
@@ 3.3 Local-Global Attention Models
${\mathbf{Y}}^{(l)}=\operatorname{Attention}^{(l)}(\text{RMS}({\mathbf{X}}^{(l)}))+{\mathbf{X}}^{(l)}$ ${\mathbf{X}}^{(l+1)}=\operatorname{SwiGLU}(\text{RMS}({\mathbf{Y}}^{(l)}))+{\mathbf{X}}^{(l)},$
@@ 4.1 Main Results
inter-chunkinter-chunkintra-chunkintra-chunk03264128256512102420480.350.350.350.350.40.40.40.40.450.450.450.45Window SizeMMLU Score (5-shot)Global Attention OnlyLocal-Global w. SWALocal-Global w. RAttentionLocal-Global w. Linear Attention
Figure 4: Comparison of MMLU 5-shot performance scores across different window sizes at 3B scale with pretraining context length 4096. The horizontal purple dashed line represents the baseline using only global attention. The blue line shows Local-Global with sliding window attention (SWA), while the red line demonstrates the performance of Local-Global with RAttention. When window size $=0$, Local-Global with RAttention reduces to Local-Global with only linear attention.

### v2

@@ Abstract
Local-global attention models characterAI2024 ; gemmateam2024gemma2 have recently emerged as compelling alternatives to standard Transformers, promising improvements in both training and inference efficiency. However, the crucial choice of window size presents a Pareto tradeoff: larger windows maintain performance akin to full attention but offer minimal efficiency gains in short-context scenarios, while smaller windows can lead to performance degradation. Current models, such as Gemma2 and Mistral, adopt conservative window sizes (e.g., 4096 out of an 8192 pretraining length) to preserve performance. This work investigates strategies to shift this Pareto frontier, enabling local-global models to achieve efficiency gains even in short-context regimes. Our core motivation is to address the intrinsic limitation of local attention—its complete disregard for tokens outside the defined window. We explore RAttention, a variant of local attention integrated with a specialized linear attention mechanism designed to capture information from these out-of-window tokens. Pretraining experiments at the 3B and 12B scales demonstrate that RAttention achieves a superior Pareto tradeoff between performance and efficiency. As a sweetspot, RAttention with a window size of just 512 consistently matches the performance of full-attention models across diverse settings. Furthermore, the recurrent nature inherent in the linear attention component of RAttention contributes to enhanced long-context performance, as validated on the RULER benchmark hsieh2024ruler . Crucially, these improvements do not compromise training efficiency; thanks to a specialized kernel implementation and the reduced window size, RAttention maintains training speeds comparable to existing state-of-the-art approaches.  [^1] Our JAX implementation of RAttention using Pallas kernels are open-sourced at https://github.com/apple/axlearn/tree/main/axlearn/common/rattention.
@@ 3.1 Residual Linear Attention (RLA)
where instead of reading out from ${\mathbf{S}}_{t}$ as in Eq 1, ${\bm{o}}_{t}$ is obtained via reading out from ${\mathbf{S}}_{t-w-1}$ – the hidden states that captures the contextual information ends at token (tw1)(t-w-1). To integrate with SWA, we combine he outputs from RLA and SWA, as demonstrated in Figure 1.
@@ 3.2 Hybrid RAttention Models
inter-chunk inter-chunk intra-chunk intra-chunk $\mathbf{S}_{[i-1]}$ ${\mathbf{S}}_{[i]}$ $\ldots$ ${\mathbf{S}}_{[\scriptscriptstyle i+m-1]}$ ${\mathbf{S}}_{[i+m]}$ ${\mathbf{O}}^{\text{inter}}_{[i]}$ ${\mathbf{O}}^{\text{intra}}_{[i]}$ ${\mathbf{O}}^{\text{inter}}_{[i+m]}$ ${\mathbf{O}}^{\text{intra}}_{[i+m]}$ ${\mathbf{Q}}_{[i]}$ ${\mathbf{K}}_{[i]}$ ${\mathbf{V}}_{[i]}$ ${\mathbf{Q}}_{[\scriptscriptstyle i+m]}$ ${\mathbf{K}}_{[\scriptscriptstyle i+m]}$ ${\mathbf{V}}_{[\scriptscriptstyle i+m]}$ Load from HBM Store to HBM On-chip $x$ Sequential
@@ 3.3 Local-Global Attention Models
${\mathbf{Y}}^{(l)}=\operatorname{Attention}^{(l)}(\text{RMS}({\mathbf{X}}^{(l)}))+{\mathbf{X}}^{(l)}\$${\mathbf{X}}^{(l+1)}=\operatorname{SwiGLU}(\text{RMS}({\mathbf{Y}}^{(l)}))+{\mathbf{X}}^{(l)},$
@@ 4.1 Main Results
03264128256512102420480.350.350.40.40.450.45Window SizeMMLU Score (5-shot)Global Attention OnlyLocal-Global w. SWALocal-Global w. RAttentionLocal-Global w. Linear Attention
Figure 4: Comparison of MMLU 5-shot performance scores across different window sizes at 3B scale with pretraining context length 4096. The horizontal purple dashed line represents the baseline using only global attention. The blue line shows Local-Global with sliding window attention (SWA), while the red line demonstrates the performance of Local-Global with RAttention. When window size =0=0, Local-Global with RAttention reduces to Local-Global with only linear attention.
@@ 5 Conclusion and Future Work
Acknowledgment
We thank Sam Wiseman, Tao Lei, Aonan Zhang, Jianyu Wang, Karen Yang for their valuable feedback.